OpenStudy (anonymous):

A rectangular prism has a width of 92 ft and a volume of 240 ft2. Find the volume of a similar prism with a width of 23 ft. Round to the nearest tenth, if necessary. (1 point) 3.8 ft3 60 ft3 15 ft3 10.4 ft3

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

@mckayela123 I answered this in your other thread :D

Directrix (directrix):

The cube of the scale factor of two similar rectangular solids is equal to the ratio of the volumes of the two similar solids. The scale factor in this case is the ratio of the widths: 92/23. Following the theorem, (92/23) ^ 3 = 240/V where V is the volume of the rectangular solid with width 23.

Directrix (directrix):

@mckayela123 Your task is to solve this equation for V. (92/23) ^ 3 = 240/V Post your work and somebody will come by to check. Hope this helps.